OpenStudy (anonymous):

((3^n)^4)((9)^(5n-2))

hartnn (hartnn):

so, \(\large (3^n)^4 \times 9^{5n-2}\) yo need to simplify this?

OpenStudy (anonymous):

yes

hartnn (hartnn):

9 can be written as ? \(\large 9=3^{??}\)

hartnn (hartnn):

for the 1st term, you can use the property that \((a^b)^c= a^{bc}\) so, \((3^n)^4=...?\)

OpenStudy (anonymous):

3^4n?

OpenStudy (anonymous):

3^(4n) x 9^(5n-2) ? is it necessary if i multiply 3 and 9? and add their exponents?

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

oh... its not so the final answer is 3^(4n) 9^(5n-2) thanks!

hartnn (hartnn):

is it necessary if i multiply 3 and 9? and add their exponents? <---incorrect 3^(4n) is correct now, \(9= 3^{??}\)

OpenStudy (anonymous):

9 = 3^2

OpenStudy (anonymous):

3^(10n-4)

OpenStudy (anonymous):

3^(6n-4)?

hartnn (hartnn):

yes, \(\large 9^{5n-2} = 3^{2 [5n-2]} = 3^{10n-4} \) got this ? now you can compare and equate the exponents of 3 :)

hartnn (hartnn):

on left its 4n on right its 10n-4 so, solve for 'n' in 4n = 10n-4 n=...?

OpenStudy (anonymous):

2/3

hartnn (hartnn):

correct! :)

hartnn (hartnn):

oh wait! you need not solve it, just simplify :P

OpenStudy (anonymous):

so = 3^(8/3) - 3^(8/3) =0

hartnn (hartnn):

so, \(\large 3^{4n} 3^{10n-4} = 3^ {4n+10n-4}= 3^{14n-4}\) thats simplified...